# Limit Theorems For Non-Hermitian Ensembles

### v3

@@ 2. Limit theorems for the complex Ginibre ensemble
2.1. Limiting left and right tail distributions of the minimum modulus as $N$ goes to infinity
@@
Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity
More precisely, for $0<r\ll 1$,
@@
The statement of Theorem 3 is derived from the formulation of the gap probability $P(r_{min}^{(N)}(A)\geq r)$ which corresponds to the $N$-th partial product presented in Equation (10).
$\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=e^{-r^{2}}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\prod_{k=1}^{N-1}e^{-r^{2}}e_{k}(r^{2})$
$=e^{-r^{2}}\prod_{k=1}^{N-1}\left[1-e^{-r^{2}}\sum_{t=n}^{+\infty}\frac{r^{2t}}{t!}\right]=e^{-r^{2}}H^{(N)}(r,0)$
where $e_{n}(x)=\sum_{k=0}^{n}\frac{x^{k}}{k!}$ defines the $n$-th partial sum of the exponential function and $H^{(N)}(r,0)$ is the conditional probability (cf. Appendix B) that given one eigenvalue lies at the origin of the complex plane all the others are found outside the disk centred at zero with radius $r$.
The limit
@@
This conditional probability $H^{(N)}(r,0)$ is a result presented in [13] and [15], where, in the limit as $N$ goes to infinity and $r$ is small (or $r$ close to zero)
@@
The limit presented in Equation (20) is derived as follows. For small $x$, the lower incomplete gamma function has asymptotic expansion
@@
$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=\prod_{k=1}^{+\infty}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}-\frac{r^{2(k+1)}}{\Gamma(k+1)}\sum_{n=1}^{+\infty}(-1)^{n}\frac{r^{2n}}{n!(k+n+1)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}+O\left(r^{2(k+2)}\right)\right]$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$
$=\exp\left[\log\left(\prod_{k=1}^{+\infty}\left[1-\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right]\right)\right]=\exp\left[\sum_{k=1}^{+\infty}\log\left(1-\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)\right]$
$=\exp\left[\sum_{k=1}^{+\infty}\left[-\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]\right]=\exp\left[-\sum_{k=1}^{+\infty}\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]$
$=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]^{j}$
@@
$\lim_{N\rightarrow+\infty}H^{(N)}(0,r)$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$
$=\left[\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}\right]$
$=1-\left[\frac{r^{4}}{2}+\frac{r^{6}}{6}+\frac{r^{8}}{24}+O(r^{10})\right]=1-O\left(r^{4}\right)$
This implies that
$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)$ $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}$
@@
Remark 2.1.
Using successive integrations by parts
$\int_{r^{2}}^{+\infty}e^{-t}t^{k}dt$ $=e^{-r^{2}}\sum_{j=0}^{k}\frac{k!}{(k-j)!}r^{2(k-j)}+k!\int_{r^{2}}^{+\infty}e^{-t}dt$ (22)
Consequently,
$P(r_{min}^{(N)}(A)\geq r)$ $=\prod_{k=0}^{N-1}\left[\frac{1}{\Gamma(k+1)}e^{-r^{2}}\sum_{j=0}^{k}\frac{k!}{(k-j)!}r^{2(k-j)}+\int_{r^{2}}^{+\infty}e^{-t}dt\right]$ (23)
@@
Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. The right tail distribution of the minimum of moduli $r_{min}^{(N)}(A)$ converges to the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{\frac{1}{k}}$ as $N$ goes to infinity, i.e., for large $r$,
@@
Let $r=\sqrt{N}|\lambda|$ such that $|\lambda|\leq 1$, where $\lambda$ is a complex number lying in the unit disk centred at the origin of the complex plane.
@@
This result corresponds to the distribution of rare events, i.e., the distribution of extreme events corresponding to the scaled minimum of the moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ as $N$ goes to infinity. This result refers to the large deviations theory. This limit at the logarithmic scale, as $N$ goes to infinity, is
@@
2.2. Independence of the scaled spectral radius and the scaled minimum modulus for the complex Ginibre ensemble
@@
Let A denote a $N\times N$ complex Ginibre matrix. The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ are independent random variables.
@@
Applying the Andreief’s integration formula
@@
From the results presented in [2], for small $r$ of order $\frac{1}{\sqrt{N}}$ (i.e. where $Nr^{2}$ is kept fixed while $N$ goes to infinity), as $N$ goes infinity (or $N$ very large) also known as the origin limit
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}E_{0}^{(2)}\left(\sqrt{N}r\right)$
Defined in [2], $E_{0}^{(\beta)}\left(\sqrt{N}r\right)$ is the probability that zero eigenvalue lies inside the disk of radius $\sqrt{N}r$ centred at the origin of the complex plane, and all eigenvalues lie outside. The index $\beta=2$ for the complex Ginibre ensemble. This implies that
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq r\right)=\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\right)$
@@
3. Limit theorems for the complex induced Ginibre ensemble
The limiting distribution of the spectral radius of non-Hermitian ensembles as well as its precise localisation near the edge of the unit disk is investigated in the work of B. Rider [20]. He dedicated his studies to the complex and symplectic Ginibre ensembles introduced by J. Ginibre [10]. These statistical ensembles share similar features such as the universality conjecture known as the circular law. Their eigenvalues move towards the unit disk as the size of the matrices increases. The following results are derived for the complex induced Ginibre ensemble using the methodological approach exposed in Rider’s work. Let $G$ denote a $N\times N$ random matrix from the complex induced Ginibre ensemble with rectangularity index $L$ defined in [5]. At appropriate scalings, the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{L+N}}$ evolves near the outer radius $r_{out}$ while the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{L}}$ is a random variable fluctuating around the inner radius $r_{in}$. The outer and inner radii are equal to $\sqrt{L+N}$ and $\sqrt{L}$, respectively.
3.1. Limit theorems of the outer and inner edges of the ring
@@
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius. Then,
@@
This limiting probability distribution is the standard Gumbel distribution for maxima. The corresponding cumulative density function is the function $F_{X}^{(\text{Gumbel max})}(x)=\exp(-\exp(-x))$.
The scaled spectral radius $R_{N}$ is approximated as
$R_{N}\approx 1+T_{\alpha,N}+\xi_{\alpha,N}$ (28)
@@
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel (maximum) distribution.
@@
This result is similar to the one detailed in [20] for the complex Ginibre ensemble. The exact formula of the probability density function $p_{R_{N}}\left(r\right)$ of the scaled spectral radius $R_{N}$, for a proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, is derived from equation (26) and is defined as follows for any value of NN
$p_{R_{N}}\left(r\right)=2[(1+\alpha)N]^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-(1+\alpha)Nr^{2}}\prod_{k=1}^{N}\frac{\gamma\left(k+\alpha N,(1+\alpha)Nr^{2}\right)}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{[(1+\alpha)N]^{j}r^{2j}}{\gamma\left(j+\alpha N,(1+\alpha)Nr^{2}\right)}\right]$ (29)
with $\gamma(.,.)$ is the lower incomplete Gamma function.
@@
The proof of Theorem 6 is presented in Appendix A.1. The analytical formulation of the probability density function $p_{R_{N}}$ of the scaled spectral radius $R_{N}$ does fit exactly its empirical distribution created from the generation of 1010 000000 complex induced Ginibre matrices with proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$. This is presented in Figure 1. As the size of these matrices increases to large numbers, the distribution of $R_{N}$ narrowing close to the outer radius $r_{out}=1$ (Figure 2).
(a) $\alpha=\frac{1}{9}$
(b) $\alpha=\frac{1}{90}(red),\alpha=\frac{1}{9}(blue)\text{ and }\alpha=\frac{4}{9}(black)$
@@
Figure 1. (left panel) Empirical probability distribution (histogram) of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=90$, $\alpha=\frac{1}{9}$ and a proportional rectangularity index $L=\alpha N$. The corresponding exact (analytical) probability density function $p_{R_{N}}$ (solid curve). (right panel) The analytical (exact) probability density function of scaled spectral radii $R_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=90$. The rectangularity index $L$ is proportional to $N$ such that $L=\alpha N$ with $\alpha>0$. The results are presented for different values of $\alpha=\{\frac{1}{90},\frac{1}{9},\frac{4}{9}\}$. Graphs generated with MATLAB.
Copyright Olivia V. Auster for code and graphs.
@@
Figure 2. Empirical probability distribution (histogram) of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=90$, $\alpha=\frac{1}{9}$ and a proportional rectangularity index $L=\alpha N$. The exact (analytical) probability density function $p_{R_{N}}$ (blue curve). Limiting probability distributions of the scaled spectral radius $R_{N}$ presented with the red and black curves for large $N=10e3$ and $N=2\times 10e3$, respectively. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
The exact formula of the cumulative density function of $R_{N}$ converges towards the asymptotic distribution as $N$ goes to infinity (Figure 3).
@@
Figure 3. (left panel) The exact cdf of the scaled spectral radius $R_{N}$ for $N=10e3$ (blue), $N=10e4$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$. The asymptotic cdf of $R_{N}$ with asymptotic location and scale parameters $\mu$ and $\sigma$, respectively (red dotted curve). (right panel) The exact formulation of the scaled spectral radius $R_{N}$ cumulative distribution (black curve) for $N=10e7$. The asymptotic cumulative distribution (red curve) parametrised with asymptotic location $\mu$ and scale $\sigma$ parameters. The parameter $\alpha=1$ for the two curves. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
@@
The scaled minimum modulus of eigenvalues, for matrices from the complex induced Ginibre ensemble with a proportional rectangularity index, follows a Gumbel (minimum) distribution at the edge of the inner circle of the ring. At the inner edge of the ring, setting $a=1-\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ is an increasing function in both $x$ and $N$, it is found that
@@
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}$ denote the scaled minimum modulus such that $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
@@
in which $\gamma_{\alpha,N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima. The scaled minimum of moduli $r_{N}$ is approximated with the following
$r_{N}\approx 1-T_{\alpha,N}+\xi_{\alpha,N}$ (32)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ where $X$ follows a standard Gumbel (minimum) distribution.
The exact formula of the probability density function $p_{r_{N}}\left(r\right)$ of the scaled minimum modulus $r_{N}$ for proportional rectangularity index $L=\alpha N$ is derived from equation (30) and is defined as follows for any value of NN

### v4

@@ 2. Methods
Asymptotic results are then derived from these analytical expressions in the limit as the size of matrices goes to infinity. They are presented in Section  3.
The limiting distribution of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for $N\times N$ complex Ginibre matrices is established in the research paper from B. Rider [22]. His framework aims to derive the logarithm of the cumulative distribution function of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for the complex Ginibre ensemble, in the limit as $N$ goes to infinity. The limiting distribution of the extreme moduli $r_{min}^{(N)}(G)$ and $r_{max}^{(N)}(G)$ and their independence for $N\times N$ matrices $G$ from the complex induced Ginibre ensemble, are derived at appropriate scaling with this framework whose main steps are detailed as follows. First, analytical formula of the survival (or cumulative) distribution function of these random variables are derived as a partial product of the cumulative distribution function of an averaged sum of independent and identically distributed random variables, each following a standard exponential distribution. As an illustration, B. Rider [22] considered the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and established that
where the random variable $Z^{(k)}=\sum_{j=1}^{k+1}Z_{j}$ follows a Gamma distribution with shape parameter $k+1$ and rate parameter $1$. The random variables $Z_{j}$, with $j\in\{1,...,k+1\}$, are independent and identically distributed. Each of the $Z_{j}$ follows a standard exponential distribution. Then, lower and upper bounds are derived for this analytical formula in the limit as $N$ goes to infinity. The limit, as $N$ goes to infinity, of the logarithm of this distribution function is then established by applying the classical Edgeworth expansion. Then, an approximation from this expression considering the leading order sum as the Riemann integral is derived. Finally, an appropriate definition of a random function is chosen such that the limiting distribution of the considered random variable, here the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$, is the identified statistical extreme distribution, here the Gumbel distribution.
3. Results
3.1. Limit theorems for the complex Ginibre ensemble
Limiting distributions of extreme moduli of matrices from non-Hermitian ensembles are of interest in the field. The limiting distribution of the spectral radius is studied in references [22, 23, 24]. Extreme moduli limiting left and right tail distribution as well as their mutual independence have not been investigated in the literature and are the themes of next Sections 3.1.1, 3.1.2, 3.2.1 and 3.2.2 for the complex Ginibre ensemble and its matrix ensemble generalisation, the complex induced Ginibre ensemble. Limiting stochastic expression of these extreme random variables are established for the complex induced Ginibre ensemble.
3.1.1. Limiting left and right tail distributions of the minimum modulus for the complex Ginibre ensemble as $N$ goes to infinity
@@ 3.1. Limit theorems for the complex Ginibre ensemble
Let $A$ denote an $N\times N$ matrix from the complex Ginibre ensemble. The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$, as $N$ goes to infinity. More precisely, for $0<r\ll 1$,
@@
The statement of Theorem 3 is derived from the formulation of the gap probability $P(r_{min}^{(N)}(A)\geq r)$ which corresponds to the $N$-th partial product presented in Equation (10),
$\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=e^{-r^{2}}\prod_{k=1}^{N-1}e^{-r^{2}}e_{k}(r^{2})=e^{-r^{2}}H^{(N)}(r,0)$
where $e_{n}(x)=\sum_{k=0}^{n}\frac{x^{k}}{k!}$ defines the $n$-th partial sum of the exponential function and $H^{(N)}(r,0)$ is the conditional probability that given one eigenvalue lies at the origin of the complex plane all the others are found outside the disk centred at zero with radius rr. [^1] The derivation of the conditional probability $H^{(N)}(r,0)$ is presented in the appendix.
Thus,
@@
The conditional probability $H^{(N)}(r,0)$ is a result presented in [14] and [17], where, in the limit as $N$ goes to infinity and $r$ is small (or $r$ close to zero)
@@
The limit presented in Equation (20) is derived as follows. The lower incomplete gamma function has series expansion, for $x$ around zero,
@@
$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}+O\left(r^{2(k+2)}\right)\right]$
@@
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=\exp\left[\sum_{k=1}^{+\infty}\left[-\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]\right]$
@@
$\lim_{N\rightarrow+\infty}H^{(N)}(0,r)=1-O\left(r^{4}\right)$
@@
Let $A$ denote an $N\times N$ matrix from the complex Ginibre ensemble. The right tail distribution of the minimum of moduli $r_{min}^{(N)}(A)$ converges to the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{\frac{1}{k}}$ as $N$ goes to infinity, i.e., for large $r$,
@@
Let $r=\sqrt{N}|\lambda|$ such that $|\lambda|\leq 1$, where $\lambda$ is a complex number lying in the unit disk centred at the origin of the complex plane. The limit
@@
This result corresponds to the distribution of extreme events of the scaled minimum modulus $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ as $N$ goes to infinity. This result refers to the large deviations theory, where
@@
3.1.2. Independence of the scaled spectral radius and the scaled minimum modulus for the complex Ginibre ensemble
@@
Let A denote an $N\times N$ complex Ginibre matrix. The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ are independent random variables, in the limit as $N$ goes to infinity.
@@
Applying Andreief’s integration formula [8],
@@
From the results presented in [2], for small $r$ of order $\frac{1}{\sqrt{N}}$ (i.e., where $Nr^{2}$ is kept fixed while $N$ goes to infinity), as $N$ goes infinity (or $N$ very large) also known as the origin limit
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}E_{0}^{(2,N)}\left(r\right)$
Defined in [2], $E_{0}^{(\beta,N)}\left(r\right)$ is the probability that zero ($\sqrt{N}\text{ scaled}$-) eigenvalue lies inside the disk of radius $r$ centred at the origin of the complex plane, and all eigenvalues lie outside. The index $\beta=2$ for the complex Ginibre ensemble. This implies that
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\right)$
@@
3.2. Limit theorems for the complex induced Ginibre ensemble
The limiting distribution of the spectral radius of non-Hermitian ensembles as well as its precise localisation near the edge of the unit disk is investigated in the work from B. Rider [22]. He dedicated his studies to the complex and symplectic Ginibre ensembles introduced by J. Ginibre [11]. These statistical ensembles share similar features such as the universality conjecture known as the circular law. Their eigenvalues (scaled with respect to $\sqrt{N}$) move towards the unit disk as the size of the matrices increases. The following results are derived for the complex induced Ginibre ensemble using the methodological approach exposed in Rider’s work.
Let $G$ denote an $N\times N$ random matrix from the complex induced Ginibre ensemble with rectangularity index $L$ as defined in [6]. The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{L+N}}$ evolves near the outer radius $r_{out}$ while the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{L}}$ is a random variable fluctuating around the inner radius $r_{in}$.
3.2.1. Limit theorems at the outer and inner edges of the ring
@@
Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius.
Then,
@@
The scaled spectral radius $R_{N}$ is approximated, as $N$ goes to infinity, by
$R_{N}\simeq 1+T_{\alpha,N}+\xi_{\alpha,N}$ (26)
@@
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel distribution for maxima.
@@
This result is similar to the one detailed in [22] for the complex Ginibre ensemble. The exact formula of the probability density function $p_{R_{N}}\left(r\right)$ of the scaled spectral radius $R_{N}$, for a proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, is derived from equation (24) and is defined as follows for any value of $N$,
$p_{R_{N}}\left(r\right)=C^{(\alpha,N)}(r)\prod_{k=1}^{N}\frac{\gamma\left(k+\alpha N,(1+\alpha)Nr^{2}\right)}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{[(1+\alpha)N]^{j}r^{2j}}{\gamma\left(j+\alpha N,(1+\alpha)Nr^{2}\right)}\right]$ (27)
where $\gamma(.,.)$ is the lower incomplete Gamma function and the function
$C^{(\alpha,N)}(r)=2[(1+\alpha)N]^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-(1+\alpha)Nr^{2}}$ (28)
@@
Numerical illustration of these results are depicted in Figures 1, 2 and 3. The analytical formulation of the probability density function $p_{R_{N}}$ of the scaled spectral radius $R_{N}$ does fit exactly its empirical distribution created from the generation of 1010 000000 complex induced Ginibre matrices with proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, as it is presented in Figure 1. As the size of these matrices increases to large numbers, the distribution of $R_{N}$ does narrow close to the outer radius $r_{out}=1$ (Figure 2).
@@
Figure 1. (left panel) Empirical probability distribution (histogram) of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=90$, $\alpha=\frac{1}{9}$ and a proportional rectangularity index $L=\alpha N$. The corresponding exact (analytical) probability density function $p_{R_{N}}$ (solid curve). (right panel) The analytical (exact) probability density function of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=90$. The rectangularity index $L$ is proportional to NN such that $L=\alpha N$ with $\alpha>0$. The results are presented for different values of $\alpha=\{\frac{1}{90},\frac{1}{9},\frac{4}{9}\}$.
@@
Figure 2. Empirical probability distribution (histogram) of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=90$, $\alpha=\frac{1}{9}$ and a proportional rectangularity index $L=\alpha N$. The exact (analytical) probability density function $p_{R_{N}}$ (blue curve). Limiting probability distributions of the scaled spectral radius $R_{N}$ presented with the red and black curves for large $N=10e3$ and $N=2\times 10e3$, respectively.
@@
Figure 3. (left panel) The exact cdf of the scaled spectral radius $R_{N}$ for $N=10e3$ (blue), $N=10e4$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$. The asymptotic cdf of $R_{N}$ (red dotted curve). (right panel) The exact formulation of the scaled spectral radius $R_{N}$ cumulative distribution (black curve) for $N=10e7$. The asymptotic cumulative distribution (red curve). The parameter $\alpha=1$ for the two curves.
@@
The scaled minimum modulus of eigenvalues, for matrices from the complex induced Ginibre ensemble with a proportional rectangularity index, follows a Gumbel (minimum) distribution at the edge of the inner circle of the ring (Theorem 7). Setting $a=1-\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ is an increasing function in both $x$ and $N$, it is found that
@@
where the $Z_{j}$ for $j=\{1,\cdots,k+L\}$ are independent and identically distributed random variables, each following a standard exponential distribution.
@@
Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ denote the scaled minimum modulus.
Then,
@@
in which $\gamma_{\alpha,N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$. This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima.
The scaled minimum of moduli $r_{N}$ is approximated as $N$ goes to infinity, by
$r_{N}\simeq 1-T_{\alpha,N}+\xi_{\alpha,N}$ (31)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$, where $X$ follows a standard Gumbel (minimum) distribution.
The exact formula of the probability density function $p_{r_{N}}\left(r\right)$ of the scaled minimum modulus $r_{N}$ for proportional rectangularity index $L=\alpha N,\forall\alpha>0,$ is derived from equation (29) and is defined as follows for any value of NN